OpenStudy (anonymous):

Porsha forgets her last homework question at school, but she is able to recall most of the problem. She needs to find the average of 5 numbers. Four of the numbers are 54, 58, 36, and 60. She remembers that the average is 12 less than the fifth number, but does not remember the fifth number or the average. Which statement best explains whether Porsha has enough information to determine the fifth number and average?

OpenStudy (anonymous):

a. Porsha does not have enough information. She needs the fifth number to determine the average. b. Porsha does not have enough information. She needs the sum of the 5 numbers to determine the average. c. Porsha does have enough information. She could solve the equation 208 + x = 5(x –12) to determine, x, the average. She can then use the average to find the fifth number. d. Porsha does have enough information. She could solve the equation 208 + x = 5(x – 12) to determine, x, the fifth number. She can then use the fifth number to determine the average.

OpenStudy (theeric):

Hi! have you read through all of the options?

OpenStudy (theeric):

My general suggestion is to look to the choices that say it \(\it is\) possible, and test to see how it works!